A container (with a movable piston) contains 20.0 L of compressed nitrogen gas at 10.0 bar and 25.0 oC. Calculate q, w, ∆ U and ∆H as one lets the gas expand reversibly and isothermally to a pressure of 1.00 bar.
1
Expert's answer
2020-02-03T08:33:24-0500


For isothermally process


"T=const"


or


"p_1V_1=p_2V_2=const"


"\Delta U=\nu C_V \Delta T=0"


"W=\frac{m}{M}RT \ln \frac{p_1}{p_2}"


"p_1V_1=\frac{m}{M}RT"


So,


"W=\frac{m}{M}RT \ln \frac{p_1}{p_2}=p_1V_1\ln \frac{p_1}{p_2}=1000000 \cdot0.02\cdot\ln \frac{1000000}{100000}\approx46 kJ"


"q=\Delta U+W=0+46=46 kJ"


"H=U+pV"


"H_1=U_1+p_1V_1"

"H_2=U_2+p_2V_2"


"\Delta H=\Delta U+p_2V_2-p_1V_1=0+p_2V_2-p_1V_1=p_2V_2-p_1V_1=0"


Answer. "W=46 kJ,q=46kJ,\Delta U=0,\Delta H=0."